John is 1/4 as old as his son.In 8 years' time,the sum of their ages will be 61. How old is John's father?
Answer by stanbon(75887)

John is 1/4 as old as his son.In 8 years' time,the sum of their ages will be 61. How old is John's father?
-------
Note: I don't know how john can be younger than his son.
==========
Equations:
j = (1/4)s
j+8 + s+8 = 61
-------
Note: s = 4j
-------
Substitute for "s" and solve for "j":
j+8 + 4j+8 = 61
5j + 16 = 61
5j = 45
john = 9 years
------
Solve for "s":
s = 4j
son = 4*9 = 36 years
--------------------------
Pls. check the wording of your problem.
John cannot be younger than his son.
=============================
Cheers,
Stan H.